OpenStudy (user123):

2x^2-5x-3=0 please factor

rishavraj (rishavraj):

Do u know about Splitting the middle term?

OpenStudy (user123):

what do you mean?

OpenStudy (will.h):

\[2x^2 - 5x - 3 = 0\] that's better.. DDo you know about groups factoring?

rishavraj (rishavraj):

@user123 http://www.purplemath.com/modules/factquad.htm have a look :)

OpenStudy (user123):

2 times 3 is 6

OpenStudy (will.h):

clearly no so what you need to do is to find 2 number with sum equal to -5 and when multiplied they must equal to -6 Any ideas about those 2 numbers?

rishavraj (rishavraj):

just a HINT - 3 - 2 = -5 -3 * (-2) = 6

OpenStudy (will.h):

Correct rishavraj well team work So those 2 numbers - 3 and -2 you need to add them instead of -5x so we would have the following \[(2x^2 - 2x )(- 3x -3 )= 0\]

OpenStudy (will.h):

actually @rishavraj is wrong he picked the wrong numbers when multiplied they must equal -6 not 6 and the sum must equal to -5 So those numbers would be -6 and 1 so therefor \[(2x^2 + x) (-6x - 3) = 0\]

rishavraj (rishavraj):

@Will.H oopps my bad...i didnt notice tht tiny "-"

rishavraj (rishavraj):

the combination would be -6 + 1 :P lol

OpenStudy (will.h):

and now we have to find the GCF for the 1st binomial which is x so therefor \[x(2x + 1)\] and now find the GCF of the 2nd binomial which is -3 so therefor \[-3(2x + 1)\] And now we have the new expression as \[(2x + 1)( x - 3)\]

OpenStudy (will.h):

yeah @rishavraj lol it is fine.. So yeah if you have any questions about this tag me

rishavraj (rishavraj):

@user123 \[2x^2 - 5x - 3 = 0\] \[2x^2 + x - 6x - 3 = 0\] \[x(2x + 1) - 3 (2x + 1) = 0\] \[(x - 3)(2x + 1) = 0\]

OpenStudy (user123):

thanks. how should you solve this with the zero product property?

OpenStudy (user123):

@Will.H

OpenStudy (will.h):

you mean you wanna find the zeros?

OpenStudy (user123):

yea it says to solve it